OpenStudy (anonymous):

Use ohm's law to solve: V=2, I=.5, Solve for R. A: 2 B: 1.5 C: 4 D: 2.5

OpenStudy (liamschumm):

Ohm's law states that v=i*r.

OpenStudy (liamschumm):

Do you get it now?

OpenStudy (anonymous):

No -.-

OpenStudy (liamschumm):

Ok, no problem. Plugging in your values gives: 2= 0.5 * R.

OpenStudy (anonymous):

Yea but what is R? -.-

OpenStudy (liamschumm):

Just solve for it using algebra. :)

OpenStudy (liamschumm):

If you want to know what it means in the formula, it means resistance.

OpenStudy (anonymous):

okay so for. Use Ohm's law to solve: I=12, R=9, Solve for V. would it be 108?

OpenStudy (liamschumm):

v= 12 * 9 = 108. Good work!

OpenStudy (anonymous):

Thanks haha and math is my best subject but with all this science stuff in here it throws me off :p

OpenStudy (anonymous):

so Use ohm's law to solve: V=30, R=6, Solve for I ; is 5?

OpenStudy (anonymous):

just want to make sure im right

OpenStudy (liamschumm):

Once you get the hang of physics, it really is mostly math.

OpenStudy (liamschumm):

Yes. 6*5=30.

OpenStudy (anonymous):

Okay just wanted to make sure that i was doing the equation correctly

OpenStudy (anonymous):

Thank you

OpenStudy (liamschumm):

You are very welcome!

OpenStudy (anonymous):

@liamschumm What is the equation for this type or problem? A 12V battery runs a 30 amp car stereo. What is the resistance of the circuit?